OpenStudy (anonymous):

Help Please! I'm about to have a meltdown You are given two complex numbers: z = 2-4i and w = 1+3i. Find the following. Give the answers in the form a+bi. Note: if a number does not have a real part, enter "0" for a. Similarly, if a number does not have an imaginary part, enter "0" for b. a) The sum: b) The difference z-w: c) The product: d) The quotient z \w: e) Simplify i^5:

OpenStudy (jtvatsim):

OK, the sum of two complex numbers is easy: (a +bi ) + (c + di) = (a+c) + (b+d)i just add the reals, and add the imaginary's

OpenStudy (jtvatsim):

same idea for the difference

OpenStudy (jtvatsim):

So can you figure out what the sum ought to be?

OpenStudy (anonymous):

yes i think so.

OpenStudy (jtvatsim):

OK, let me know what you get. :)

OpenStudy (anonymous):

3-7i ?

OpenStudy (anonymous):

am i distributing?

OpenStudy (jtvatsim):

(2 - 4i) + (1 + 3i) = 2 - 4i + 1 + 3i = 2 + 1 - 4i + 3i = 3 - i

OpenStudy (jtvatsim):

treat the i's like x's in algebra and you will get the right answer

OpenStudy (anonymous):

thanks!